https://sites.google.com/a/pessto.org/wiki/pessto-wiki/home/night-reports/night-report-2018-02-05

Night Report 2018-02-05

Observers: Alessandro Razza, Regis Cartier
Support Team: Owen McBrien, Peter Clark, Erkki Kankare
TAT:  Janet Chen, Michel Dennefeld
TO: Javier Alarcon (day), Mauricio Martinez (night)

Calibrations:  EFOSC PESSTO_calib, PESSTO_bias, PESSTO_calib_gr18+gr20.

Weather: Thin clouds/Cirrus for the majority of the night. Telescope closed owing to thick clouds at UT 7:28.

UT            Object                setup        exp        airmass    seeing        note
---------------------------------------------------------------------------------------------------------------------------------------------------------------
EFOSC
00:56:10    SN2017imj     gr13 sl1    2700s        2.20        1.0
01:54:10    Focus                                                                                    TO had to focus again
02:04:50    A2018hu        gr13 sl1    900s          1.55        0.86            Classification --> Type II z=0.051. Strong host galaxy contamination
02:28:01    OGLE17jei     gr11 sl1    2700s        1.61        0.66
03:20:10    OGLE17jei     gr16 sl1    1800s        1.75        0.66             Problem with Flats
04:02:55    OGLE17jei     gr16 sl1                                                           FLATS ONLY
04:24:10    LTT3864        gr11+13+16                1.09        0.76
04:40:33    AT2018ph     gr13 sl1    300s           1.25        0.72            Classification --> Ia around a week before max. z=0.03
04:52:32    SN2018oh     gr11+16    900s          1.51        0.72            Flats ok
SOFI 
05:52:49    Hip44089                                          1.88        0.59            Telluric
06:06:21    SN2018oh    BG            2160s         1.64        0.59            Only two spectra before thick clouds. We did not abort the OB.
06:25:00                                                                                                   Thick Clouds
07:10:03                                                                                                   Telescope Downtime
07:28:00                                                                                                   Telescope Closed                                                                   
       
Unfortunately, critical infrared observations could not be completed due to thick clouds. We plan to observe with EFOSC tomorrow.


Comments